# ETHIC: Evaluating Large Language Models on Long-Context Tasks with High Information Coverage

### v1

@@ Abstract
Recent advancements in large language models (LLM) capable of processing extremely long texts highlight the need for a dedicated evaluation benchmark to assess their long-context capabilities. However, existing methods, like the needle-in-a-haystack test, do not effectively assess whether these models fully utilize contextual information, raising concerns about the reliability of current evaluation techniques. To thoroughly examine the effectiveness of existing benchmarks, we introduce a new metric called information coverage (IC), which quantifies the proportion of the input context necessary for answering queries. Our findings indicate that current benchmarks exhibit low IC; although the input context may be extensive, the actual usable context is often limited. To address this, we present ETHIC, a novel benchmark designed to assess LLMs’ ability to leverage the entire context. Our benchmark comprises 2,648 test instances spanning four long-context tasks with high IC scores in the domains of books, debates, medicine, and law. Our evaluations reveal significant performance drops in contemporary LLMs, highlighting a critical challenge in managing long contexts. Our benchmark is available at https://github.com/dmis-lab/ETHIC.
@@
Taewhoo Lee1  Chanwoong Yoon1  Kyochul Jang2  Donghyeon Lee1  Minju Song1 Hyunjae Kim1,†Jaewoo Kang1,3,† 1Korea University  2Seoul National University  3AIGEN Sciences {taewhoo,cwyoon99,dong9733,thdalswn99}@korea.ac.kr kyochul@snu.ac.kr{hyunjae-kim,kangj}@korea.ac.kr
@@ 2.3 Information Coverage
A few concurrent studies have also sought to establish the criteria for evaluating long-context LLMs. Goldman et al. (2024) defined the aspect of “scope” as “how much necessary information is there to find?,” which is similar to our definition. However, they did not provide a specific metric and systematically evaluate existing benchmarks. In contrast, we present a novel approach for measuring the quantity of information, marking the first time this has been done in this area.
@@ 3 ETHIC
This section provides details on the dataset and task construction process.
@@ 3.1 Corpus Selection
Table 2: A summary of our dataset construction.  ETHIC covers four domains—books, debates, medicine, and law—with a total of 2,648 test instances. In the medicine and law domains (marked with $\dagger$), the inputs consist of multiple documents, while the inputs in the books and debates domains consist of a single long document. We randomly sampled 50 instances per domain and task to report information coverage.
@@ 3.3 Annotation Process
Since annotating long contexts all at once can reduce accuracy, we processed each context by dividing it into smaller chunks (up to 1,024 tokens). We used GPT-4o to annotate each small chunk, and merged the labels without duplicates to obtain the final label set for the full input document. We initially instructed GPT-4o to review its own answers, but empirically found that this process often led the model to misjudge correct answers as incorrect. We manually reviewed the accuracy of the model’s annotations by examining 100 chunk-label pairs from each domain and found that they were highly accurate (see Section E for details).
@@ 4 Experiments
Table 3: The performance of models and training-free methods on ETHIC. The best scores are highlighted in bold. We used instruction-tuned versions of open-source models. Refer to Section 4.1 for the details of the metrics.
@@ 5 Analysis
We conducted further analysis to better understand the models’ limitations within our benchmark. We used GPT-4o mini in the experiments.
@@ 5.1 Comparison of Model Performance Between Low-IC and High-IC Tasks
As illustrated in Figure 1, we previously observed that model performance can vary significantly based on the amount of information required to answer the query, even when using the same input contexts. To explore this further, we constructed a set of low-IC tasks, including traditional NLP tasks such as single-document QA, multi-document QA, and query-focused summarization (QFS). Specifically, we extracted one or more chunks from the contexts in our benchmark and asked GPT-4 to generate queries and their corresponding answers. This method followed a framework similar to that used for developing our benchmark, ensuring high annotation accuracy. Figure 3 shows the performance was consistently better on low-IC tasks compared to high-IC tasks. This highlights that, even when tasks involve the same documents and similar cognitive demands, information coverage is a key factor that significantly impacts model performance.
Figure 3: The model’s performance on low-IC and high-IC tasks. Low-IC tasks were created by generating new queries and answers using the same input context from our benchmark, which are represented by the gray bars on the left side of the graph (refer to Section 5.1 for details). The yellow bars on the right represent high-IC tasks from our benchmark. The numbers (%) displayed in the bar graphs represent the IC values of the tasks. The y-axis indicates the model performance.
@@ 5.3 Effect of Position of Information
A recent study demonstrated that when relevant information is located in the middle of the input sequence, both QA and retrieval performance significantly declined Liu et al. (2024). To investigate whether a similar position-dependent phenomenon occurs in our tasks, we visualized the model scores for each chunk in the summarizing task. The figure shows that the model effectively processed information at the beginning of the provided context, but performance decreased toward the end. For context lengths ranging from 8K to 32K, the performance in the middle sections was the lowest, which is consistent with findings from Liu et al. (2024). When the context length exceeds 32K, we consistently observed a decline in performance as the context length increased. This might occur because, during decoding, the summaries of the later chunks are influenced by the outputs from the earlier sections, which increases the likelihood of errors in those later parts. However, this does not imply that the errors in the later sections are solely due to decoding issues. The performance of the earlier sections also decreased as the context length increased, suggesting that the model had inherent limitations in its ability to handle long-context encoding.
Figure 5: The effect of the position of information in the summarizing task. The x-axis indicates the position of the chunks within the input context, while the y-axis represents the total length of the input context.
@@ 6 Conclusion
We calculated the IC value for long contexts by breaking them into smaller chunks and assessing whether each chunk is necessary to answer the given query. However, a limitation of this approach is that, in cases requiring multi-hop reasoning, some chunks may mislead the evaluator into believing they do not directly contribute to answering the query, even if they are crucial for intermediate reasoning when connected to other chunks. Consequently, the evaluator might overlook the potential usefulness of these chunks when evaluating them individually. Our study did not thoroughly analyze this aspect. In future research, we will conduct additional analyses and improvements in this area.
@@ Acknowledgements
This research was supported by (1) the National Research Foundation of Korea (NRF2023R1A2C3004176, RS-2023-00262002), (2) the Ministry of Health & Welfare, Republic of Korea (HR20C0021(3)), and (3) ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2024-20200-01819).
@@ Appendix A Details on Information Coverage
We outline the prompt used to measure IC values as follows:
You are given a passage and a query. The given query originally requires a reader to answer based on a longer context. This task splits the long context into multiple small passages, and aims to identify whether a certain passage needs to be taken into account in order to answer the query. ### Passage: (Note: This passage is a small portion of the original context, and may not be provided in the format specified in the query.) <context> ### Query: (Note: This query is for evaluation only. DO NOT answer it yourself.) <query> ### Score Rubrics: 0 : The passage is an unnecessary part of the original context, which does not need to be taken into account to answer the query. 1 : The passage is a necessary part of the original context, which should be taken into account to answer the query. Follow the following format: # Query Understanding : {{demonstrate a clear understanding of the query and its requirements}} # Passage Understanding : {{demonstrate a clear understanding of the passage, focusing on how it relates to the query}} # Assessment : {{demonstrate your assessment based on the Score Rubrics provided above}} # Final Score : {{a single score only}}
@@ Appendix B Details on Annotation Process
To evaluate performance, we constructed prompts using dataset chunks (<passages>). In the case of the recall task, the same prompts were also used to construct the dataset. The prompts for each dataset by task are as follows.
B.1 Recall
@@ Appendix E Validation of Data Quality
We labeled a subset of the recalling task dataset manually for quality control. The F1 scores comparing GPT4o-generated labels and human annotations are presented in Table B.

### v2

@@ Abstract
Recent advancements in large language models (LLM) capable of processing extremely long texts highlight the need for a dedicated evaluation benchmark to assess their long-context capabilities. However, existing methods, like the needle-in-a-haystack test, do not effectively assess whether these models fully utilize contextual information, raising concerns about the reliability of current evaluation techniques. To thoroughly examine the effectiveness of existing benchmarks, we introduce a new metric called information coverage (IC), which quantifies the proportion of the input context necessary for answering queries. Our findings indicate that current benchmarks exhibit low IC; although the input context may be extensive, the actual usable context is often limited. To address this, we present ETHIC, a novel benchmark designed to assess LLMs’ ability to leverage the entire context. Our benchmark comprises 1,986 test instances spanning four long-context tasks with high IC scores in the domains of books, debates, medicine, and law. Our evaluations reveal significant performance drops in contemporary LLMs, highlighting a critical challenge in managing long contexts. Our benchmark is available at https://github.com/dmis-lab/ETHIC.
@@
Taewhoo Lee1,3  Chanwoong Yoon1  Kyochul Jang2  Donghyeon Lee1,3  Minju Song1 Hyunjae Kim1,†Jaewoo Kang1,3,† 1Korea University  2Seoul National University  3AIGEN Sciences {taewhoo,cwyoon99,dong9733,thdalswn99}@korea.ac.kr kyochul@snu.ac.kr{hyunjae-kim,kangj}@korea.ac.kr
@@ 2.3 Information Coverage
A few concurrent studies have also sought to establish the criteria for evaluating long-context LLMs. Goldman et al. (2024) defined the aspect of “scope” as “how much necessary information is there to find?”, which is similar to our definition. However, they did not provide a specific metric and systematically evaluate existing benchmarks. In contrast, we present a novel approach for measuring the quantity of information, marking the first time this has been done in this area.
@@ 3 ETHIC
In this section, we outline the document collection process (see Section 3.1) and the task construction process (see Section 3.2) in detail. Additionally, we describe the label annotation process for each task in Section 3.3.
@@ 3.1 Corpus Selection
Table 2: A summary of our dataset construction.  ETHIC covers four domains—books, debates, medicine, and law—with a total of 1,986 test instances. In the medicine and law domains (marked with $\dagger$), the inputs consist of multiple documents, while the inputs in the books and debates domains consist of a single long document. We randomly sampled 50 instances per domain and task to report information coverage.
@@ 3.3 Annotation Process
Since annotating long contexts all at once can reduce accuracy, we processed each context by dividing it into smaller chunks (up to 1,024 tokens). We used GPT-4o to annotate each small chunk, and merged the labels without duplicates to obtain the final label set for the full input document. We initially instructed GPT-4o to review its own answers, but empirically found that this process often led the model to misjudge correct answers as incorrect. We manually reviewed the accuracy of the model’s annotations by examining 100 chunk-label pairs from each domain and found that they were highly accurate (see Appendix E for details).
@@ 4 Experiments
Table 3: The performance of models and training-free methods on ETHIC. The best scores are highlighted in bold. We used instruction-tuned versions of open-source models. Please refer to Section 4.1 for the details of the metrics.
In this section, we outline different metrics used to evaluate each task (see Section 4.1). We introduce baseline models and methods in Section 4.2, and provide the results in Section 4.3.
@@ 5 Analysis
We conducted further analysis to better understand the models’ limitations within our benchmark. We used GPT-4o mini as the main model throughout Sections 5.1 to 5.3, and all models for Section 5.4.
@@ 5.1 Comparison of Model Performance Between Low-IC and High-IC Tasks
As illustrated in Figure 1, we observed that model performance can vary significantly based on the amount of information required to answer the query, even when using the same input contexts. To explore this further, we constructed a set of low-IC tasks, including traditional NLP tasks such as single-document QA, multi-document QA, and query-focused summarization (QFS). Specifically, we extracted one or more chunks from the contexts in our benchmark and asked GPT-4o to generate queries and their corresponding answers. This method followed a framework similar to that used for developing our benchmark, ensuring high annotation accuracy. Figure 3 shows the performance was consistently better on low-IC tasks compared to high-IC tasks. This highlights that, even when tasks involve the same documents and similar cognitive demands, information coverage is a key factor that significantly impacts model performance.
Figure 3: The model’s performance on low-IC and high-IC tasks. Low-IC tasks were created by generating new queries and answers using the same input context from our benchmark, which are represented by the gray bars on the left side of the graph (please refer to Section 5.1 for details). The yellow bars on the right represent high-IC tasks from our benchmark. The numbers (%) displayed in the bar graphs represent the IC values of the tasks. The y-axis indicates the model performance.
@@ 5.3 Effect of Position of Information
A recent study demonstrated that when relevant information is located in the middle of the input sequence, both QA and retrieval performance significantly declined Liu et al. (2024). To investigate whether a similar position-dependent phenomenon occurs in our tasks, we visualized the model scores for each chunk in the summarizing task. Figure 5 shows that the model effectively processed information at the beginning of the provided context, but performance decreased toward the end. For context lengths ranging from 8K to 32K, the performance in the middle sections was the lowest, which is consistent with findings from Liu et al. (2024). When the context length exceeds 32K, we consistently observed a decline in performance as the context length increased. This might occur because, during decoding, the summaries of the later chunks are influenced by the outputs from the earlier sections, which increases the likelihood of errors in those later parts. However, this does not imply that the errors in the later sections are solely due to decoding issues. The performance of the earlier sections also decreased as the context length increased, suggesting that the model had inherent limitations in its ability to handle long-context encoding.
Figure 5: The effect of the position of information in the summarizing task. The x-axis represents the position of the chunks within the input context, while the y-axis represents the total length of the input context. Blue (red) chunks indicate summaries with high (low) scores.
5.4 Error Analysis: Degeneration
Through a close examination of model responses, we noticed that models mainly suffer from degeneration Welleck et al. (2020); Holtzman et al. (2020); Fu et al. (2021); Li et al. (2023), i.e. generation of unreasonably repetitive texts. Following previous works, we used Rep-n and Rep-r to quantify this phenomenon across different models (see Appendix F for details of the metrics). Specifically, we applied Rep-n on the Recalling task to account for the portion of repeated n-grams, and Rep-r on the Summarizing task to account for the portion of repetitive summary snippets. As shown in Table 4, all models exhibited severe degeneration issues across both tasks. This indicates that tasks requiring high information coverage adversely impacted model behavior during generation, ultimately hindering performance. Furthermore, notably high repetition scores observed from models such as GPT-4o indicates that degeneration remains a critical issue even for bigger models. Exploring different strategies to mitigate degeneration would be crucial for improving model performance on tasks that demand high information coverage.
Model Recalling Summarizing
Rep-n (%) Rep-r (%)
Proprietary
Gemini Pro 1.5 27.9 43.0
GPT-4o 90.9 65.2
GPT-4o mini 69.8 53.8
Open-Source
Qwen2.5-72B-Instruct 87.3 61.0
Llama-3.1-70B-Instruct 70.7 67.7
Qwen2.5-7B-Instruct 82.9 65.5
Llama-3.1-8B-Instruct 76.8 66.0
GLM4-9B-Chat 89.9 63.6
Phi-3.5-mini-instruct 55.4 58.3
Table 4: Rep-n and Rep-r scores across different models on the Recalling and Summarizing tasks. Please refer to Appendix F for the details of the metrics.
@@ 6 Conclusion
We calculated the IC value for long contexts by breaking them into smaller chunks and assessing whether each chunk is necessary to answer the given query. However, in cases requiring multi-hop reasoning, some chunks may mislead the evaluator into believing they do not directly contribute to answering the query, even if they are crucial for intermediate reasoning when connected to other chunks. Consequently, the evaluator might overlook the potential usefulness of these chunks when evaluating them individually. While our study did not thoroughly analyze this aspect, we encourage future research to conduct additional analyses and improvements in this area.
@@ Acknowledgements
This research was supported by (1) the National Research Foundation of Korea (NRF2023R1A2C3004176, RS-2023-00262002), (2) the Ministry of Health & Welfare, Republic of Korea (HR20C0021), (3) ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2025-20200-01819), and (4) Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency(KOCCA) grant funded by the Ministry of Culture, Sports and Tourism(MCST) in 2023(Project Name: Development of storytelling AI technology for cultural heritage tailored to the various interests of users, Project Number: RS-2023-00220195, Contribution Rate: 100%).
@@ Appendix A Details on Information Coverage
Below is the instruction used to measure IC.
You are given a passage and a query. The given query originally requires a reader to answer based on a longer context. This task splits the long context into multiple small passages, and aims to identify whether a certain passage needs to be taken into account in order to answer the query. ### Passage: (Note: This passage is a small portion of the original context, and may not be provided in the format specified in the query.) <passages> ### Query: (Note: This query is for evaluation only. DO NOT answer it yourself.) <query> ### Score Rubrics: 0 : The passage is an unnecessary part of the original context, which does not need to be taken into account to answer the query. 1 : The passage is a necessary part of the original context, which should be taken into account to answer the query. Follow the following format: # Query Understanding : {{demonstrate a clear understanding of the query and its requirements}} # Passage Understanding : {{demonstrate a clear understanding of the passage, focusing on how it relates to the query}} # Assessment : {{demonstrate your assessment based on the Score Rubrics provided above}} # Final Score : {{a single score only}}
@@ Appendix B Details on Annotation Process
Below are the instructions used to evaluate models across different tasks. Note that for the Recalling task, the same instructions were used in the annotation process, but with smaller chunks of text.
B.1 Recalling
@@ Appendix E Validation of Data Quality
To ensure the quality of labels generated by GPT-4o, we randomly selected 100 samples from each domain. The resulting annotations were evaluated against GPT-4o-generated labels using the F1 score, as shown in Table B.
@@
Appendix F Details on Degeneration Analysis
We used two metrics to measure degeneration: Rep-n and Rep-r. Following the notations from Li et al. (2023), each metric is calculated as follows:
Rep-n $=1.0-\frac{|\text{UniqueNgrams}(x,n)|}{L-n+1}$
[…]
$\vee(x_{i}=x_{k}\wedge x_{i-1}=x_{k-1},\exists k\neq i)\Big{\}}\Big{|}$
where $x$, $L$, and n refers to the sentence, its length, and the length of n-gram within the sentence, respectively. Rep-n measures repetition based on the portion of repeated n-grams, whereas Rep-r quantifies repetition based on the portion of repeated snippets measured by sentence length.